OpenStudy (anonymous):

If I have sqrt56/8 can I simplify that to sqrt7?

OpenStudy (anonymous):

\[\sqrt{56}/8 = \sqrt{7} or no?

OpenStudy (anonymous):

what would it be? if it isnt

OpenStudy (anonymous):

\[\sqrt{\frac{ 56 }{ 8 }}\] Is this the problem?

OpenStudy (anonymous):

If so, you have the correct answer. Do you understand why?

OpenStudy (anonymous):

no this is the problem \[\frac{\sqrt{56}}{8} \]

OpenStudy (anonymous):

Ok. Try to simplify the top and see what happens

OpenStudy (anonymous):

\[\frac{ 2\sqrt{14} }{ 8 }\] is that correct?

OpenStudy (anonymous):

That is right

OpenStudy (anonymous):

thank you! i have a whole sheet of these and I was just really confused

OpenStudy (anonymous):

Don't forget to simplify more

OpenStudy (anonymous):

huh? more

OpenStudy (anonymous):

Looks like you know what you are doing

OpenStudy (anonymous):

simplify more though?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

You can simplify the fraction a bit more

OpenStudy (anonymous):

to what? I dont understand how this:? \[\frac{ \sqrt{14} }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 8 } \sqrt{14}\] This is equivalent to what you have

OpenStudy (anonymous):

Yes you got it!

OpenStudy (anonymous):

yay! THANK U!

OpenStudy (anonymous):

You are welcome